OpenStudy (anonymous):

The surface area of a right circular cylinder of height 5 feet and radius r feet is given by S(r)=2πrh+2πr2. Find the instantaneous rate of change of the surface area with respect to the radius, r, when r = 6.

zepdrix (zepdrix):

\[\Large\rm S(r)=2\pi r h+2\pi r^2\]Hey Camila! :) Have you learned your derivative shortcuts at this point? Or do we have to use the limit definition for the derivative to find our instantaneous rate of change? -_-

OpenStudy (anonymous):

I know how to derive.

zepdrix (zepdrix):

We're holding the height, h, constant at 5.\[\Large\rm S(r)=2\pi r \cdot 5+2\pi r^2\]Simplifies our function a little bit, which is nice.\[\Large\rm S(r)=10\pi r+2\pi r^2\]

zepdrix (zepdrix):

So apply your power rule to each term, should be pretty straight forward :) Are you getting confused because of all that pi nonsense mixed in?

OpenStudy (anonymous):

Yeah that was the part that confused me.

zepdrix (zepdrix):

\[\Large\rm (10\pi x)'=10\pi(x)'=10\pi (1)\]Normal power rule, ignoring the constant coefficients :)

OpenStudy (anonymous):

So I would get 10π+4πr and plug in 6 for r?

zepdrix (zepdrix):

\[\Large\rm S'(r)=10 \pi+4\pi r\]Good good good.

OpenStudy (anonymous):

So when I plug that in it would be 34π right?

zepdrix (zepdrix):

Yayyy good job! *\c:/* Camila <*c:/*

OpenStudy (anonymous):

Haha okay thank you so much!

zepdrix (zepdrix):

np